A series RLC circuit withR = 6 ohm, C = 0.02 Farad and L = 0.1 has no applied voltage. Find the subsequent current in the circuit if the initial charge, on the capacitor is q0 and the initial current is zero.
1
Expert's answer
2020-03-16T13:16:02-0400

Task:

"R = 6 \Omega; \\
L = 0.1 \text{H}; \\
C = 0.02 \text{F}; \\
i(0) = 0 \text{A}; \\
q_C(0) = q_0"

Find: "i(t)"

Solution:



Current through the capacitor and voltage across the inductance can be described as:

"i_{C}=C \cfrac{d v_{C}}{d t}\ (1)\\
v_{L}=L \cfrac{d i_{L}}{d t}\ (2)"

In this case all the elements are connected in series, so "i_C = i_L = i" . According to the second Kirchhoff’s circuit law:

"v_{R}+v_{L}+v_{C}=0\ (3)"

After inserting (1) and (2) into (3) the last equation can be rewritten as:

"v_R + v_L + v_C = Ri + L \cfrac{di}{dt} + \cfrac{1}{C} \int\limits_0^{t_0} idt = 0\ (4)"

After differentiation of the left and right parts of equation (4) it is possible to get:

"\cfrac{d}{dt} (Ri + L \cfrac{di}{dt} + \cfrac{1}{C} \int\limits_0^{t_0} idt) = \cfrac{d0}{dt} \Leftrightarrow R \cfrac{di}{dt} + L \cfrac{d^2 i}{dt} + \cfrac{1}{C} i = 0.1\cfrac{d^2 i}{dt} + 6\cfrac{di}{dt} + 50i=0"

It is a second order ordianary differential equation with constant coefficients. To solve this equation it is necessary to find the roots of the characteristic polynomial:

"0.1 \lambda^2 + 6 \lambda + 50 = 0"

According to the Vieta's formuals:

"\begin{cases}
\lambda_1+\lambda_2 = -\cfrac{6}{0.1} = -60 \\
\lambda_1 \cdot \lambda_2 = \cfrac{50}{0.1} = 500
\end{cases} \Bigg| \Rightarrow 
\begin{cases}
\lambda_1 = -50 \\
\lambda_2 = -10
\end{cases}"

So, the general solution for "i(t)" is equal to:

"i(t) = C_1 e^{\lambda_1 t} + C_2 e^{\lambda_2 t} = C_1 e^{-50t} + C_2 e^{-10t}\ (5)"

To get the partial solution it is necessary to solve the Cauchy's problem. First of all its is necessary to get initial condition for "\cfrac{di}{dt}" . According to (4):

"L \cfrac{di}{dt} = -Ri-u_C \Rightarrow \cfrac{di}{dt} (0) = \cfrac{1}{L} (-Ri(0)-u_C(0)) \ (6)"

To get "u_C(0)" it is convenient to use the defenition of capacity:

"C = \cfrac{q_C}{u_C} \Rightarrow u_C(0) = \cfrac{q_C(0)}{C} = \cfrac{q_0}{C}\ (7)"

After inserting (7) into (6) the latter can be rewritten as:

"\cfrac{di}{dt}(0) = \cfrac{1}{L} (-Ri(0) - \cfrac{q_0}{C})\ (8)"

According to the task, "i(0) = 0", so, according to (8):

"\cfrac{di}{dt}(0) = -\cfrac{1}{LC} q_0 = -500q_0"

Now it is possible to solve the Cauchy's problem:

"\begin{cases}
i(0) = C_1 e^{-50\cdot 0} + C_2 e^{-10 \cdot 0} = C_1 + C_2 = 0 \\
\cfrac{di}{dt} (0) = \cfrac{d}{dt} (C_1 e^{-50t} + C_2 e^{-10t})(0) =-50C_1 -10C_2 =-500q_0
\end{cases}"

So, the system for "C_1" and "C_2" is:

"\begin{cases}
C_1+C_2=0\\
-50C_1-10C_2=-500 q_0
\end{cases} \Leftrightarrow 
\begin{cases}
C_2 = -C_1\\
-40C_1=-500 q_0
\end{cases} \Leftrightarrow 
\begin{cases}
C_1 = 12.5 q_0\\
C_2 = -12.5q_0
\end{cases} \ (9)"

After inserting (9) in (5) it is possible to get partial solution:

"i(t) = 12.5q_0 e^{-50t} -12.5q_0 e^{-10t}"

Answer: "i(t) = 12.5q_0 e^{-50t} -12.5q_0 e^{-10t}" .